OpenStudy (trewolf):

Raul is building a house. On day 1, he uses 3 bags of cement. On each of the following 3 days, he uses 3 times the number of bags used the previous day. How many bags of cement does he use in the last 3 days?

satellite73 (satellite73):

how many days does he work?

OpenStudy (.sammi.):

I think it is 3 days

satellite73 (satellite73):

it says "last three days" not "first three days"

OpenStudy (trewolf):

6?

satellite73 (satellite73):

idk it doesn't say

OpenStudy (.sammi.):

That's true...

OpenStudy (trewolf):

this is 6th grade math...

satellite73 (satellite73):

if it is the first three days then \[3+3\times 3+3\times 3\times 3\]

OpenStudy (trewolf):

Do 3+3 First!

satellite73 (satellite73):

but it says "each of the following 3 days' so maybe it is \[3\times 3+3\times 3\times 3+3\times 3\times 3\times 3\]

satellite73 (satellite73):

or if you prefer \[3^2+3^3+3^4\]

OpenStudy (trewolf):

189?